If |y+8|+2=6,then y+8=+/-4 Is this true or false? If false,please correct |y+8| +2=6.
Thanks for taking the time to help me.

Answer by addingup(3677)

|y+8|+2 = 6
= y+8 = 4 or y+8 = -4
y = -4 or y = -12
:
There's your answer.
John